I have a word problem that says: A rectangular scrapbook page has an area of 180 inches squared. If the length of the page is two less than twice the width, find the width of the page.
Answer by checkley77(12844)

L*W=180
L=2W-2
(2W-2)W=180
2W^2-2W-180=0
2(W^2-W-90)=0
2(W-10)(W+9)=0
W-10=0
W=10 ANS.
L=2*10-2
L=20-2
L=18 ANS.
PROOF
18*10=180
180=180